# Chemical Cartography with APOGEE: Two-process Parameters and Residual Abundances for 288,789 Stars from Data Release 17

### v1

@@ 6 Summary
Overall, older open clusters show smaller median residual abundances, i.e., they are more chemically similar to MW disk stars. The two youngest clusters show noticeable enhancement in median C+N, Na, S, and Cu compared to older clusters. Ce enhancement is generally seen in clusters with ages up to 2 Gyr.
@@ Appendix A Calibration Offsets
All calibration offsets, including the grid of values and values of are provided in a table, available online at https://zenodo.org/records/10659205. The first five rows are shown in Table 2. We also include a Jupyter notebook with sample code to load the table, extract the offset grid for a specific element, and obtain calibrated abundances for ten example stars. Also included in the notebook is code to reproduce Figure 4. Note that the reported calibration offsets are derived for and thus applicable to only the data used in this work (described in detail in Section 2.1).
@@
Note. — Full table available online at https://zenodo.org/records/10659205.
@@ Appendix C Catalog User Guide
The residual abundance catalog is available online at https://zenodo.org/records/10659205. The table columns are reported in Table 5. We also include a Jupyter notebook with sample code to cross-match with external catalogs, using AstroNN (Leung & Bovy, 2019a, b) and the DR17 globular cluster catalog (Schiavon et al., 2023) as examples. The notebook also includes code to make some plots, including reproducing the residual abundances (right side) of Figure 15 and plotting as function of position.

### v2

@@ 6 Summary
Overall, older open clusters show smaller median residual abundances, i.e., they are more chemically similar to MW disk stars. The two youngest clusters show noticeable enhancement in median C+N, Na, S, and Cu compared to older clusters. Ce enhancement is generally seen in clusters with ages up to 2 Gyr, similarly to young field stars. The Ce and C+N enhancements are likely a signature of enhanced AGB enrichment.
@@ Appendix A Calibration Offsets
All calibration offsets, including the grid of values and values of are provided in a table, available online at https://doi.org/10.5281/zenodo.10659205. The first five rows are shown in Table 2. We also include a Jupyter notebook with sample code to load the table, extract the offset grid for a specific element, and obtain calibrated abundances for ten example stars. Also included in the notebook is code to reproduce Figure 4. Note that the reported calibration offsets are derived for and thus applicable to only the data used in this work (described in detail in Section 2.1).
@@
Note. — Full table available online at https://doi.org/10.5281/zenodo.10659205.
@@ Appendix C Catalog User Guide
The residual abundance catalog is available online at https://doi.org/10.5281/zenodo.10659205. The table columns are reported in Table 5. We also include a Jupyter notebook with sample code to cross-match with external catalogs, using AstroNN (Leung & Bovy, 2019a, b) and the DR17 globular cluster catalog (Schiavon et al., 2023) as examples. The notebook also includes code to make some plots, including reproducing the residual abundances (right side) of Figure 15 and plotting as function of position.